### v2

Abstract
The form-factor bootstrap is incomplete until one normalizes the zero-particle form factor. For the stress energy tensor we describe how to obtain the vacuum energy density ${\rho_{\rm vac}}$, defined as $\langle 0|T_{\mu\nu}|0\rangle=\rho_{\rm vac}\,g_{\mu\nu}$, from the form-factor bootstrap. Even for integrable QFT’s in D=2 spacetime dimensions, this prescription is new, although it reproduces previously known results obtained in a different and more difficult thermodynamic Bethe ansatz computation. We propose a version of this prescription in D=4 dimensions. For these even dimensions, the vacuum energy density has the universal form $\rho_{\rm vac}\propto m^{D}/\mathfrak{g}$ where $\mathfrak{g}$ is a dimensionless interaction coupling constant which can be determined from the high energy behavior of the S-matrix. In the limit $\mathfrak{g}\to 0$, $\rho_{\rm vac}$ diverges due to well understood UV divergences in free quantum field theories. If we assume the the observed Cosmological Constant originates from the vacuum energy density $\rho_{\rm vac}$ computed as proposed here, then this suggests there must exist a particle which does not obtain its mass from spontaneous symmetry breaking in the electro-weak sector, which we designate as the “zeron”. A strong candidate for the zeron is a massive Majorana neutrino.
I Introduction
Bootstrap ideas, which originated in the 1960’s as an attempt to understand the strong interactions, have found great success in D=2 spacetime dimensions, in particular for conformal field theory (CFT) BPZ and integrable massive theories ZamoZamo0 . In recent years the bootstrap has been developed in some detail in higher dimension D with surprising success. For CFT’s see the review Poland . More recently the bootstrap has been studied for massive theories Paulos1 ; Homrich ; Karateev and the latter is more relevant to the present work.
In this article we are primarily concerned with the vacuum energy density ${\rho_{\rm vac}}$ defined as a vacuum expectation value of the stress-energy tensor $T_{\mu\nu}$:
$\langle 0|T_{\mu\nu}|0\rangle={\rho_{\rm vac}}\,g_{\mu\nu},$ (1)
with the convention $\{g_{\mu\nu}\}={\rm diag}\{1,-1,-1,-1,\ldots\}$. From this definition of the vacuum energy density ${\rho_{\rm vac}}$ one has
${\rho_{\rm vac}}=\langle 0|\Theta|0\rangle/D,~{}~{}~{}~{}~{}\Theta=T_{\mu}^{\mu}.$ (2)
We thus focus on form factors for the trace of the stress energy tensor $\Theta$. For CFT’s, $\Theta=0$. The form factor bootstrap in principle relates $n$-particle form factors to those with $n-2$ particles, thus we address the problem of determining ${\rho_{\rm vac}}$ from the 2-particle form factor. For integrable theories in $D=2$ the form-factor bootstrap is very well developed Smirnov with many applications MussardoBook . In any dimension, even in $2D$, the form factor bootstrap for any operator ${\cal O}$ is incomplete since the basic equations are linear in the form factor and are invariant under a rescaling by an arbitrary constant. Thus the form factors for any operator ${\cal O}$ are incomplete until one specifies its vacuum expectation value $\langle 0|{\cal O}|0\rangle$. This necessarily comes from additional ultra-violet (UV) data such as the thermodynamic Bethe ansatz (TBA). A main result of this paper is a prescription for determining $\langle 0|\Theta|0\rangle$ from the 2-particle form factor. This prescription is new, even in $2D$, and as we will show it reproduces known results for ${\rho_{\rm vac}}$ previously obtained from the TBA, but without introducing a finite temperature. More importantly it has a natural generalization to D=4.
In ALCC a non-perturbative definition of ${\rho_{\rm vac}}$ was proposed based on thermodynamics. In the “Thermal” channel euclidean time is compactified on a circle of circumference $\beta=1/T$ where $T$ is the temperature. Thus states are the usual scattering states in infinite $d=D-1$ dimensional volume. This is in contrast to the spatially compactified channel “SpC”, where states partially live on a circle of one compactified spatial dimension. We showed that these two channels lead to the same result for free massive theories, however the Thermal channel is better behaved since most integrals are already convergent. The results of this paper are implicitly based on the Thermal channel in the limit $\beta\to 0$. Let us summarize the non-perturbative definition of ${\rho_{\rm vac}}$ in the thermal channel proposed in Mingling ; ALCC . Let $p$ denote the pressure and ${\cal F}$ the free energy density. Assuming the theory has a single mass scale $m$, such as the mass of the lightest particle, we define the scaling variable $r=m\beta$. Then one can write
$p\,\beta^{\scriptscriptstyle{D}}=-\beta^{\scriptscriptstyle{D}}{\cal F}\equiv\chi(D)\,c(r),~{}~{}~{}~{}\chi(D)\equiv\pi^{-{\scriptscriptstyle{D}}/2}\Gamma(D/2)\zeta(D).$ (3)
The function $\chi(D)$ is the famous combination discovered by Riemann that satisfies the functional equation $\chi(D)=\chi(1-D)$ which is essential for establishing Modularity for the Thermal verses SpC channels ALCC . If the theory is UV complete and a CFT then one expects
$c(r)=c_{\rm uv}+c_{\scriptscriptstyle{D}}\,r^{\scriptscriptstyle{D}}+\ldots$ (4)
where $\ldots=\sum_{\mathfrak{p}}c_{\mathfrak{p}}\,r^{\mathfrak{p}}$ refers to terms with different powers of $r$ coming from perturbation theory about the CFT by relevant operators. Then we proposed
${\rho_{\rm vac}}=-c_{\scriptscriptstyle{D}}\chi(D)m^{\scriptscriptstyle{D}}.$ (5)
The normalization of $c(r)$ is such that $c_{\rm uv}=1$ for a free massless boson for any $D$. In $D=4$ dimensions one expects the various powers $\mathfrak{p}$ to be irrational and shouldn’t mix with $c_{\scriptscriptstyle{D}}$, unlike in 2D where conformal perturbation theory can lead to divergent terms with powers $r^{2}$ due to the rationality of anomalous dimensions of the relevant perturbation about the UV CFT. In the Thermal channel, the $c_{\scriptscriptstyle{D}}$ term is difficult to calculate even with the TBA, and this provides the main motivation for this article, namely to determine ${\rho_{\rm vac}}$ without having to calculate the full free energy density ${\cal F}$ at arbitrary temperature $T$. Our prescription below can thus be viewed as extracting ${\rho_{\rm vac}}$ from a proper $\beta\to 0$, i.e. high temperature (UV) limit. For integrable quantum field theories (QFT)’s with diagonal scattering, using the TBA one can find a simple formula for ${\rho_{\rm vac}}$ originally due to Al. Zamolodchikov ZamoTBA ; KlassenMelzer
${\rho_{\rm vac}}=\dfrac{m_{1}^{2}}{2\mathfrak{g}},$ (6)
where $m_{1}$ is the lightest particle and $\mathfrak{g}$ is a finite dimensionless interaction coupling which can be extracted from the S-matrix for the scattering of $m_{1}$ with itself (see below for a review). Using D=4 Lattice QCD results for two light quarks and one massive one (the strange quark) we used (5) to estimate ${\rho_{\rm vac}}$ and obtained the reasonable value ${\rho_{\rm vac}}\approx-(200{\rm MeV})^{4}$, and this calculation relied on asymptotic freedom Wilczek ; Gross .
A main goal of this article is to obtain a formula analogous to (6) in higher dimensions directly from the S-matrix and the form factor bootstrap. Towards this the following remark is important. For free particles there is an unavoidable divergence in the UV limit $\beta\to 0$ for D even. Specifically, for free theories ($\mathfrak{g}=0$) one finds ALCC
${\rho_{\rm vac}}=\pm\begin{cases}\dfrac{m^{2}}{4\pi}\left(\log(4\pi/m\beta)+\tfrac{1}{2}-\gamma_{E}\right)~{}~{}~{}~{}~{}~{}~{}(D=2)\\ \\ \dfrac{m^{4}}{32\pi^{2}}\left(\log(4\pi/m\beta)+\tfrac{3}{4}-\gamma_{E}\right)~{}~{}~{}~{}~{}(D=4)\end{cases}$ (7)
where +/+/-+ / - corresponds to bosons/fermions and $\gamma_{E}$ is the Euler constant. Remarkably the result (6) shows that interactions, i.e. $\mathfrak{g}\neq 0$, can regularize the divergence as $\beta\to 0$ in the free theory. This implies that UV divergences in the non-interacting theory can be absorbed into physical masses $m_{\rm phys}$ in order obtain a physical ${\rho_{\rm vac}}$. The zero temperature S-matrix depends only on these physical masses and coupling $\mathfrak{g}$ and should render a finite result for ${\rho_{\rm vac}}$. A main result of this article is such a formula for $D=4$, and based on this discussion we expect that $\lim_{\mathfrak{g}\to 0}{\rho_{\rm vac}}=\infty$. Henceforth $m$ refers to these physical masses.
One aspect of our analysis is the principle of a particular democracy. Namely since the vacuum $|0\rangle$ has no quantum numbers and cannot be excited as a resonance in scattering, then in principle the scattering of any particle with it’s anti-particle can probe ${\rho_{\rm vac}}$, and they should all agree222“Every particle is created equal when it comes to experiencing the void.”. For $D=2$ integrable theories we will show how this principle follows from the S-matrix bootstrap in the specific example of Toda theories based on the affine extension of the the Lie group $SU(N+1)$.
We present our results as follows. In the next section we describe very general properties of the 2-particle form factors of the stress-energy tensor which apply to any dimension D and don’t rely on the bootstrap nor integrability. This section mainly just serves to fix the normalization of the form factors in terms of physical masses. In Section III we turn to D=2 integrable theories and propose a prescription for how to obtain ${\rho_{\rm vac}}$ from the 2-particle form factor based on the bootstrap. Although the final result for ${\rho_{\rm vac}}$ is known to be of the form (6) by other methods in particular the TBA, our derivation directly from the S-matrix without relying on the TBA is new and actually simpler. We present the example of the affine Toda theories for SU(N+1) in order to explain how the particular democracy originates from the S-matrix bootstrap. The $SU(2)$ case is the sinh-Gordon model and we carry out a detailed analysis in this case using the explicit formula for the 2-particle form-factor. In Section IV we take some steps towards higher dimensions. There we first rewrite the $D=2$ result using the Mandelstam variable $s$. This leads us to propose a formula for ${\rho_{\rm vac}}$ in $D=4$ based entirely on the high energy behavior of the S-matrix. Since ${\rho_{\rm vac}}$ is central to any discussion of the Cosmological Constant Problem, in Section V we speculate on potential implications for physics beyond the Standard Model. We argue that if ${\rho_{\rm vac}}$ is enough to explain the observed value of the cosmological constant $\rho_{\Lambda}$, then this requires at least one stable massive particle in the deep UV above spontaneous symmetry breaking of the electro-weak sector. With lack of a better name we refer to such a hypothetical particle as the “zero-on” , where “zero” refers to the vacuum $|0\rangle$, with mass $m_{\rm z}$. In order to explain the measured value of $\rho_{\Lambda}$, it turns out that a massive Majorana neutrino is a good candidate for the zeron, since it is thought that its mass cannot yet be explained by the Higgs mechanism, and its proposed mass based on phenomenology is in the right ball-park to explain the observed value of the Cosmological Constant.
II Generalities of 2-particle form factors of the stress-energy tensor
Our primary interest is the zero particle form factor for the trace of the energy momentum tensor which determines ${\rho_{\rm vac}}$ according to (2). We wish to obtain ${\rho_{\rm vac}}$ from $2$ particle form factors. In this section we present some general, i.e. model independent, properties of these $2$-particle form factors in arbitrary spacetime dimension. As stated in the Introduction, this serves to fix the normalization of form factors in terms of physical masses.
Consider a relativistic QFT in $D=d+1$ spacetime dimensions. For simplicity let us assume the theory consists of a single particle of mass $m$. One particle asymptotic scattering states are denoted as $|{\bf k}\rangle$, where ${\bf k}$ is the $d$-dimension momentum, and are on the mass shell with associated energy momentum $D$-vector $p_{\bf k}$:
$p_{\bf k}=(\omega_{\bf k},{\bf k}),~{}~{}~{}~{}~{}\omega_{\bf k}=\sqrt{{\bf k}^{2}+m^{2}},~{}~{}~{}~{}~{}p_{\bf k}^{2}=\omega_{\bf k}^{2}-{\bf k}^{2}=m^{2}.$ (8)
Since $d^{d}{\bf k}/\omega_{\bf k}$ is Lorentz invariant, we define the resolution of the identity as follows:
${\bf 1}=\sum_{n=0}^{\infty}\dfrac{1}{n!}\int\dfrac{d^{d}{\bf k}_{1}}{(2\pi)^{d}\omega_{{\bf k}_{1}}}\cdots\int\dfrac{d^{d}{\bf k}_{n}}{(2\pi)^{d}\omega_{{\bf k}_{n}}}~{}|{\bf k}_{1},{\bf k}_{2},\ldots{\bf k}_{n}\rangle\langle{\bf k}_{1},{\bf k}_{2},\ldots{\bf k}_{n}|$ (9)
where
$\langle{\bf k}^{\prime}|{\bf k}\rangle=(2\pi)^{d}\,\omega_{\bf k}\,\delta^{(d)}({\bf k}^{\prime}-{\bf k}).$ (10)
The mass dimension [^3] Henceforth all scaling dimensions refer to mass units. of the states $|{\bf k}\rangle$ is thus $(1-d)/2$. For any quantum field operator ${\cal O}(x)$ one can define its form factors
$F^{\cal O}({\bf k}_{1},{\bf k}_{2},\ldots,{\bf k}_{n})=\langle 0|{\cal O}|{\bf k}_{1},{\bf k}_{2},\ldots{\bf k}_{n}\rangle$ (11)
where in the above equation ${\cal O}={\cal O}(0)$. Labeling particle states as $|p_{\bf k}\rangle$, crossing symmetry reads
$\langle p_{{\bf k}_{1}}|{\cal O}(0)|p_{{\bf k}_{2}},\ldots p_{{\bf k}_{n}}\rangle=\langle 0|{\cal O}(0)|-p_{{\bf k}_{1}},p_{{\bf k}_{2}},\ldots p_{{\bf k}_{n}}\rangle$ (12)
Let us now consider the operator ${\cal O}=T_{\mu\nu}$, the stress energy tensor. Since $\partial^{\mu}T_{\mu\nu}=0$, one can define an operator $A(x)$ where:
$T_{\mu\nu}(x)=\left(\partial_{\mu}\partial_{\nu}-g_{\mu\nu}\partial^{2}\right)\,A(x).$ (13)
Define the trace of stress-energy tensor $\Theta\equiv T_{\mu}^{\mu}$. Then (13) implies
$\langle 0|\Theta|{\bf k}^{\prime},{\bf k}\rangle=-d\,\,\dfrac{(p_{{\bf k}^{\prime}}+p_{{\bf k}})^{2}}{({\bf k}^{\prime}+{\bf k})^{2}}~{}\langle 0|T_{00}|{\bf k}^{\prime},{\bf k}\rangle$ (14)
Using crossing symmetry:
$\langle{\bf k}^{\prime}|\Theta|{\bf k}\rangle=-d\,\,\dfrac{(p_{{\bf k}^{\prime}}-p_{{\bf k}})^{2}}{({\bf k}^{\prime}-{\bf k})^{2}}~{}\langle{\bf k}^{\prime}|T_{00}|{\bf k}\rangle.$ (15)
One can show that
$\lim_{{\bf k}^{\prime}\to{\bf k}}\dfrac{(p^{\prime}-p)^{2}}{({\bf k}^{\prime}-{\bf k})^{2}}=-1.$ (16)
Thus
$\langle{\bf k}|\Theta|{\bf k}\rangle=d\,\langle{\bf k}|T_{00}|{\bf k}\rangle.$ (17)
One also has
$\int d^{d}{\bf x}\,\langle{\bf k}^{\prime}|T_{00}(x)|{\bf k}\rangle=\int d^{d}{\bf x}\,e^{i({\bf k}^{\prime}-{\bf k})\cdot{\bf x}}\,\langle{\bf k}^{\prime}|T_{00}(0)|{\bf k}\rangle=\langle{\bf k}^{\prime}|H|{\bf k}\rangle=\omega_{k}\,\langle{\bf k}^{\prime}|{\bf k}\rangle$ (18)
where $H$ is the hamiltonian. This formally implies $\langle{\bf k}|T_{00}(0)|{\bf k}\rangle=\omega_{\bf k}^{2}$, however based on (17), since the LHS is Lorentz invariant, $\omega_{\bf k}^{2}$ should be replaced by $m^{2}$. One thus finds simply
$\langle{\bf k}|\Theta|{\bf k}\rangle=d\,m^{2},$ (19)
which is independent of ${\bf k}$. The above equation is what fixes the normalization of the $2$-particle form factors, and this normalization is implicit for all the higher $n\neq 2$ particle form factors in the bootstrap. The 1-particle form factors are also non-zero in general and by Lorentz invariance are constant:
$\langle 0|\Theta|{\bf k}\rangle={\rm constant}.$ (20)
It is important to note that in obtaining (19) we fixed by hand ${\bf k}^{\prime}={\bf k}$ under crossing symmetry in (15) based on Lorentz invariance. This kind of manipulation of limits will be necessary below. On the other hand the two-particle form factor $\langle 0|\Theta|{\bf k},{\bf k}^{\prime}\rangle$ contains a great deal more information on the underlying dynamics, and the goal is to extract ${\rho_{\rm vac}}$ from it. In the next section we propose a prescription for doing so in the case of integrable theories in $D=2$.
III ${\rho_{\rm vac}}$ from the 2-particle form factor for D=2 Integrable QFT
Let us consider an integrable QFT in $D=2$ spacetime dimensions where the particles are labeled by “a” with physical mass $m_{a}$. We assume that each particle $a$ is its own anti-particle $\overline{a}$ for simplicity, otherwise the formulas below involve charge conjugation matrices which do not affect our main results. Integrability implies no particle production so that the scattering of 2 particles leads only to 2-particle asymptotic states. Also for simplicity, we assume the S-matrix is diagonal, which is to say that scattering of particles $a,b$ only produces particles of the same type, thus one only need consider $S_{ab}=S_{ab}^{ab}$ matrix elements.
III.1 Review of form factor axioms
In this subsection we give a brief review of the main formulas we will need. For a more comprehensive review see for instance Smirnov ; MussardoBook .
The 2-particle form factors are functions of $s\equiv(p_{1}+p_{2})^{2}$ by Lorentz invariance. For D=2, analytic properties of the S-matrix are more easily described in terms of the rapidity $\theta$:
$p=(\omega_{{\bf k}},{\bf k})=(m\cosh\theta,m\sinh\theta)$ (21)
which gives
$s=(p_{a}+p_{b})^{2}=m_{a}^{2}+m_{b}^{2}+2m_{a}m_{b}\cosh\theta_{ab},~{}~{}~{}~{}~{}\theta_{ab}\equiv\theta_{a}-\theta_{b}.$ (22)
[…]
${\bf 1}=\sum_{n=0;\{a\}}^{\infty}\,\dfrac{1}{n!}\int\dfrac{d\theta_{1}}{2\pi}\cdots\int\dfrac{d\theta_{n}}{2\pi}~{}|\theta_{1},\theta_{2},\ldots\theta_{n}\rangle_{\{a_{1}a_{2}\ldots\}}\,{}_{\{a_{1},a_{2}\ldots\}}\langle\theta_{1},\theta_{2},\ldots\theta_{n}|,~{}~{}~{}~{}_{a}\langle\theta^{\prime}|\theta\rangle_{b}=\delta_{ab}2\pi\delta(\theta-\theta^{\prime}).$ (23)
For the remainder of this section $F$ refers to the form factors of $\Theta$:
$F_{a_{1}a_{2}a_{3}\ldots}(\theta_{1},\theta_{2},\theta_{3},\ldots)=\langle 0|\Theta|\theta_{1},\theta_{2},\theta_{3},\ldots\rangle_{a_{1}a_{2}a_{3}\ldots}$ (24)
Since the states $|\theta\rangle$ have zero mass dimension, all multi-particle form factors $F$ have the same mass dimension of $2$. Form factors for any operator ${\cal O}$ satisfy well-known axioms Smirnov . Although we are only interested in the 2-particle form factors, for reasons that will be clear let us review these axioms for up to $4$ particles.
Crossing symmetry.    $p\to-p$ corresponds to $\theta\to\theta+i\pi$ in (21). This implies
$F_{a_{1}a_{2}a_{3}a_{4}}(\theta_{1}+i\pi,\theta_{2},\theta_{3},\theta_{4})=~{}_{a_{1}}\langle\theta_{1}|\Theta|\theta_{2},\theta_{3},\theta_{4}\rangle_{a_{2}a_{3}a_{4}}.$ (25)
The Watson equation.
$F_{a_{1}a_{2}a_{3}a_{4}}(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=S_{a_{1}a_{2}}(\theta_{12})\,F_{a_{2}a_{1}a_{3}a_{4}}(\theta_{2},\theta_{1},\theta_{3},\theta_{4}).$ (26)
Kinematic poles. There are generic poles from the annihilation of a particle with its anti-particle which relates the $n$ particle form factor to that for $n-2$ particles:
$-i\lim_{\theta_{1}\to\theta_{2}}(\theta_{1}-\theta_{2})~{}F_{a\overline{a}a_{3}a_{4}}(\theta_{1}+i\pi,\theta_{2},\theta_{3},\theta_{4})=\left(1-S_{aa_{3}}(\theta_{23})\,S_{aa_{4}}(\theta_{24})\right)\,F_{a_{3}a_{4}}(\theta_{3},\theta_{4}).$ (27)
Bound state poles. If particle $c$ appears at a pole in $S_{ab}(\theta)$ at $\theta=iu_{ab}^{c}$ then there is a bound state with mass
$m_{c}^{2}=m_{a}^{2}+m_{b}^{2}+2m_{a}m_{b}\cos u_{ab}^{c}.$ (28)
The scattering of particle $c$ with any other particle $d$ satisfies the bootstrap equation
$S_{cd}(\theta)=S_{ad}(\theta+i\overline{u}_{ac}^{b})\,S_{bd}(\theta-i\overline{u}_{bc}^{a}),$ (29)
and the form factor also has such poles. Starting with the S-matrix for the lightest particle $m_{1}$ with itself, the bootstrap can be closed, i.e. all other $m_{a}$’s and $S_{ab}$ can be determined.
Since the above form factor axioms are linear in $F$, they are incomplete since they do not fix the normalization of the $F$’s for the operator $\Theta$. Thus additional physical input is required. In particular
$\langle 0|\Theta|0\rangle={\rho_{\rm vac}}/D,~{}~{}{\rm and}~{}~{}\langle 0|\Theta|\theta\rangle_{a}\equiv F_{a},~{}~{}~{}~{}{\rm are~{}constant}$ (30)
due to Lorentz invariance. Such input must come from the UV since the two point correlation function of the stress-energy tensor $\langle\Theta(x)\Theta(0)\rangle$ can be expressed in terms of form factors using the resolution of the identity and if the theory is UV complete this correlation function is fixed at high energies by the UV CFT. We henceforth assume the QFT is UV complete.
III.2 A prescription for ${\rho_{\rm vac}}$
Suppose we are given the 2-particle form factors $F_{ab}(\theta_{1},\theta_{2})=F_{ab}(\theta_{12})$. The kinematic pole equation (27) in general relates $n$-particle form factors the the $n-2$ ones. However this equation makes no sense for $n=2$ as written. However some version of it should be valid with some prescription, and this is what we will present.
Consider the formal crossing relation
${}_{a}\langle\theta_{1}|\Theta|\theta_{2}\rangle_{b}=F_{ab}(\theta_{1}+i\pi,\theta_{2}).$ (31)
The form factors are assumed to be normalized such that (19) is satisfied, which requires setting $\theta_{1}=\theta_{2}=\theta$ by hand and then using Lorentz invariance to conclude the result should be independent of $\theta$:
${}_{a}\langle\theta|\Theta|\theta\rangle_{a}=\lim_{\theta_{1}\to\theta_{2}}\,F_{aa}(\theta_{1}+i\pi,\theta_{2})=m_{a}^{2}.$ (32)
This properly normalizes the form factor with no additonal freedom to change it. This kind of treatment of the rapidities will play an essential role for ${\rho_{\rm vac}}$ and the one point form factors $F_{a}$ below. An important point is that if one does not directly set $\theta_{1}=\theta_{2}$ in (31), there is still a lot of dynamical information in the form factor. The second point is that one does not expect any kinematic poles as resonances for the vacuum $|0\rangle$, precisely because it is the vacuum. We will show this explicitly below for the case of the sinh-Gordon model. Thus one does not need to cancel the pole in (27). This leads us to propose the following. The order of limits is important. We first consider the high energy limit $\theta_{1}\to\infty$ first, then subsequently set $\theta_{1}=\theta_{2}$. We denote this order of limits as $\lim_{\theta_{1}\to\infty|\theta_{1}=\theta_{2}}$. We now propose the prescription:
$-i\lim_{\theta_{1}\to\infty|\theta_{1}=\theta_{2}}\,F_{a\overline{a}}(\theta_{1}+i\pi,\theta_{2})=\lim_{\theta_{1}\to\infty|\theta_{1}=\theta_{2}}\left(1-S_{a\overline{a}}(\theta_{12})\right)\,\langle 0|\Theta|0\rangle.$ (33)
The above equation is a dramatic example of the mingling between the UV and the infra-red, since the LHS is essentially fixed by masses that are measured at low energy, whereas the RHS involves the extremely high energy limit of the S-matrix.
The above formula (33) leads to a simple formula for ${\rho_{\rm vac}}$. In general the basic building blocks of the two-body S-matrices $S_{ab}(\theta)$ are factors of $f_{\alpha}(\theta)$:
$S_{ab}(\theta)=\prod_{\alpha\in{\cal A}_{ab}}f_{\alpha}(\theta),~{}~{}~{}~{}~{}f_{\alpha}(\theta)\equiv\dfrac{\sinh\tfrac{1}{2}\left(\theta+i\pi\alpha\right)}{\sinh\tfrac{1}{2}\left(\theta-i\pi\alpha\right)},$ (34)
where ${\cal A}_{ab}$ is a finite set of $\alpha$’s. One has
$\lim_{\theta\to\pm\infty}f_{\alpha}(\theta)=e^{\pm i\pi\alpha}\left(1\pm 2i\sin(\pi\alpha)\,e^{-|\theta|}+{\cal O}(e^{-2|\theta|})\right).$ (35)
Thus
[…]
where we have used $\sum_{\alpha\in{\cal A}_{ab}}\,\alpha=0$. The prescription (33) combined with (32) then leads to
$m_{a}^{2}=\mathfrak{g}_{aa}\,\langle 0|\Theta|0\rangle.$ (37)
Since the vacuum has no quantum numbers, the equation (37) must lead to the same result for $\langle 0|\Theta|0\rangle$ for any particle of type $a$. Namely any particle $a$ is able to probe $\langle 0|\Theta|0\rangle$, and this property was referred to as particular democracy in the Introduction. This is where the bootstap comes in. Using that the energy and momentum operators are conserved quantities, combined with the bootstrap equation (29), one can show: [^4] See for instance KlassenMelzer .
$\mathfrak{g}_{ab}=\widehat{m}_{a}\widehat{m}_{b}\,\,\mathfrak{g}_{11},~{}~{}~{}~{}\widehat{m}_{a}\equiv m_{a}/m_{1},$ (38)
where $\mathfrak{g}_{11}$ is for the lightest particle $m_{1}$. Then (37) leads to, for any particle $a$:
$\langle 0|\Theta|0\rangle=\dfrac{m_{1}^{2}}{\mathfrak{g}_{11}},~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\rho_{\rm vac}}=\dfrac{m_{1}^{2}}{2\mathfrak{g}_{11}}.$ (39)
The above result agrees with what is obtained from a thermodynamic approach, namely the TBA ZamoTBA . This new derivation of ${\rho_{\rm vac}}$ is considerably simpler and doesn’t depend on a finite temperature treatment. As we will see, this has its advantages, especially for its generalization to higher dimensions where a TBA treatment is not possible.
A related but slightly different limit constrains the one-particle form factors. Even for the stress-energy tensor, these one-particle form factors are constants and are not necessarily zero. For instance for the magnetic perturbation of the Ising model, $\Theta$ is the spin field $\sigma$ and has non-zero 1-particle form factors DelfinoMussardo ; DelfinoCardy ; DelfinoE8 . It was conjectured in these works that when a subset of the rapidities are taken to infinity, the form factors should factorize as a result of a kind of cluster decomposition. Let us assume this conjecture. In (33) if one only takes the $\theta_{1}\to\infty$ limit without also the subsequent limit $\theta_{1}\to\theta_{2}$ the form factors should factorize into 1-particle form factors. We thus propose:
$\lim_{\theta_{1}\to\infty}\,F_{ab}(\theta_{1},\theta_{2})=\dfrac{F_{a}\,F_{b}}{\langle 0|\Theta|0\rangle},~{}~{}~{}~{}~{}F_{a}\equiv\langle 0|\Theta|\theta\rangle_{a}~{}={\rm constant}.$ (40)
The above formula only makes sense if the LHS is finite and independent of rapidities. The latter follows from
$\lim_{\theta\to\infty}S(\theta)=1,$ (41)
which can be verified for the examples below.
III.3 Example: Affine Toda theories
The affine Toda theories provide a nice illustrative example since they have multiple particles of mass $m_{a}$ and the scattering is diagonal (for real coupling $b$ below). They also have a coupling constant $b$ such that the exact result for ${\rho_{\rm vac}}$ can be compared with perturbation theory. We consider only the simply laced Lie algebras $G$, the ADE series, in particular $G=A_{N}=SU(N+1)$. Although the final result is known from the TBA, we present it here as an illustration of the prescription (33) and particular democracy.
Denote the simple roots of $SU(N+1)$ as $\bm{\alpha}_{i}$, $i=1,2,\dots N$:
$\bm{\alpha}_{i}=(\alpha_{i}^{1},\alpha_{i}^{2},\ldots,\alpha_{i}^{N})\equiv\{\alpha_{i}^{a},~{}a=1,2,\ldots N\}.$ (42)
Introduce $N$ real scalar fields $\bm{\phi}$:
$\bm{\phi}=(\phi^{1},\phi^{2},\ldots,\phi^{N})\equiv\{\phi^{a},~{}a=1,2,\ldots,N\}$ (43)
The untwisted affine Lie algebra has one additional root $\bm{\alpha}_{0}=-\sum_{i=1}^{N}\bm{\alpha}_{i}$. With these definitions one can define the $2D$ Euclidean action
${\cal S}=\int d^{2}x\left(\dfrac{1}{8\pi}\partial\bm{\phi}\cdot\partial\bm{\phi}+V(\bm{\phi})\right),~{}~{}~{}~{}~{}V(\bm{\phi})=\mu\,\sum_{i=0}^{N}e^{b\,\bm{\alpha}_{i}\cdot\bm{\phi}}.$ (44)
The $1/8\pi$ normalization corresponds to standard $2D$ conformal field theory conventions where $\langle\phi^{a}(x)\phi^{b}(0)\rangle=-\delta_{ab}\log x^{2}$ which fixes the convention for the coupling $b$. With this convention $V(\bm{\phi})$ is a strongly relevant perturbation of anomalous scaling dimension $-2b^{2}$ if the all roots are conventionally normalized as $\alpha_{i}^{2}=2$. The masses satisfy the relation
$\widehat{m}_{a}=\dfrac{m_{a}}{m_{1}}=\dfrac{\sin(a\pi/(N+1))}{\sin(\pi/(N+1)}.$ (45)
The affine Toda theories have many applications. For real coupling $b$ the S-matrices were first obtained in Fateev . For $SU(2)$ the model is the sinh-Gordon model. The more interesting physical applications are for purely imaginary coupling $b$, where for $SU(2)$ this is the sine-Gordon model. For the latter the scattering is not diagonal except at the so-called reflectionless points. For other Lie groups $G$, the affine Toda theories have solitons in the spectrum and the scattering is not diagonal and satisfies the Yang-Baxter equation BernardLeClair ; Hollowood . The quantum theory for imaginary $b$ has a quantum affine symmetry ${\cal U}_{q}(\widehat{G})$ where $\widehat{G}$ is the affine extension of $G$ BernardLeClair . For $q$ a root of unity, some RSOS restrictions of these theories describe integrable perturbations of minimal CFT’s AhnBL . The RSOS restricted $SU(3)$ theory describes the 3-state Potts model. Restriction of the $G=E_{8}$ theory describes magnetic perturbations of the Ising model and one can reproduce the results of Zamolodchikov ZamoE8 , where the theory is diagonal under RSOS restriction. For the present article we only consider real coupling $b$ since the theory is diagonal for arbitrary $b$. A comprehensive review of such theories can be found in Braden .
The S-matrix bootstrap can be completed starting with the just the lightest mass particle of mass $m_{1}$. Let $S_{11}(\theta)$ denote the S-matrix for this particle with itself:
$S_{11}=\prod_{\alpha\in{\cal A}_{11}}f_{\alpha}=f_{\tfrac{2}{h}}\,f_{-\tfrac{2\gamma}{h}}\,f_{\tfrac{2(\gamma-1)}{h}},~{}~{}~{}h=N+1,~{}~{}~{}~{}~{}~{}\gamma\equiv\dfrac{b^{2}}{1+b^{2}}.$ (46)
($h$ is the dual Coxeter number.) Then
$\mathfrak{g}_{11}=2\sum_{\alpha\in{\cal A}_{11}}\sin\pi\alpha=-8\sin(\pi/h)\sin(\pi\gamma/h)\sin(\pi(1-\gamma)/h).$ (47)
Based on equation (39), one then finds
${\rho_{\rm vac}}=-\dfrac{m_{1}^{2}}{16\,\sin(\pi/h)\,\sin(\pi\gamma/h)\,\sin(\pi(1-\gamma)/h)}.$ (48)
Expanding ${\rho_{\rm vac}}$ for $N=1$ (the sinh-Gordon model) one finds
$\lim_{b\to 0}\,{\rho_{\rm vac}}=-m^{2}\left(\dfrac{1}{8\pi b^{2}}-\dfrac{1}{8\pi}-\dfrac{\pi b^{2}}{48}+\dfrac{\pi b^{4}}{48}-\dfrac{(60+7\pi^{2})\pi b^{6}}{2880}+\ldots\right)~{}~{}~{}~{}~{}(N=1).$ (49)
Rescaling $m\to b^{2}m$ the above defines a well-defined perturbative expansion that can be compared with Feynman diagram perturbation theory, which was understood by Destri-deVega by summing over all tadpole Feynman diagrams in a rather complicated calculation DestriDeVega .
III.4 Analytical details for the sinh-Gordon model
The detailed analysis of this section is not essential for our purposes since we obtained (39) without knowing the exact 2-particle form factor function, but it’s useful to analyze an example where the 2-particle form factors are explicitly known so as to verify some of the above properties. The affine Toda theory for $SU(2)$ is the sinh-Gordon model with the action
${\cal S}=\int d^{2}x\left(\dfrac{1}{8\pi}(\partial_{\mu}\phi\,\partial^{\mu}\phi)+2\mu\cosh(\sqrt{2}\,b\,\phi)\right).$ (50)
The spectrum consists of a single particle of physical mass $m$. From (46) one has
$S(\theta)=-f_{-\gamma}(\theta)f_{\gamma-1}(\theta)=\dfrac{\sinh\theta-i\sin\pi\gamma}{\sinh\theta+i\sin\pi\gamma},~{}~{}~{}~{}~{}\gamma\equiv b^{2}/(1+b^{2}).$ (51)
This already implies
$\mathfrak{g}=-4\sin\pi\gamma,~{}~{}~{}~{}~{}\Longrightarrow~{}~{}\langle 0|\Theta|0\rangle=-\dfrac{m^{2}}{4\sin\pi\gamma}=2{\rho_{\rm vac}}.$ (52)
and ${\rho_{\rm vac}}$ is negative.
Form factors for the sinh-Gordon model were studied in great detail by Fring, Mussardo and Simonetti FringMussardo . Henceforth $F(\theta_{1},\theta_{2})=F(\theta_{12})$ is the two particle form factor for $\Theta$. From the Watson equation we define a minimal solution $F_{\rm min}$:
$F_{\rm min}(\theta)=S(\theta)\,F_{\rm min}(-\theta).$ (53)
Form factors for ${\cal O}$ for $n>2$ numbers of particles factorize into a product of $F_{\rm min}(\theta_{ij})$ times a polynomial in $e^{\theta_{i}}$, where the polynomial depends on the operator ${\cal O}$. $F_{\rm min}$ has the integral representation [^5] For numerical evaluation of this integral, the hybrid formula eq. 4.36 in FringMussardo is very useful. See also Negro for some more recent analysis of $F_{\rm min}$.:
$\log F_{\rm min}(\theta)=-4\int_{0}^{\infty}\dfrac{dx}{x}~{}\dfrac{\sinh(x\gamma/2)\sinh(x(1-\gamma)/2)\sinh(x/2)}{\sinh^{2}(x)}\,\cdot\cos(\widetilde{\theta}x/\pi),~{}~{}~{}~{}~{}~{}~{}~{}\widetilde{\theta}\equiv i\pi-\theta.$ (54)
Note that $F_{\rm min}(\theta+i\pi)$ is real if $\theta$ is real. The properly normalized form factor is
$F(\theta_{12})=\langle 0|\Theta|\theta_{1},\theta_{2}\rangle=m^{2}\,\dfrac{F_{\rm min}(\theta_{12})}{F_{\rm min}(i\pi)}.$ (55)
With this normalization
$\langle\theta|\Theta|\theta\rangle=\lim_{\theta\to 0}\dfrac{F_{\rm min}(\theta+i\pi)}{F_{\rm min}(i\pi)}=m^{2}.$ (56)
Let us first show that there is no kinematic resonance poles in $F(\theta_{1}+i\pi,\theta_{2})=F(\theta_{12}+i\pi)$, as anticipated above, such that the residue axiom (27) makes no sense for the 2-particle form factor. For this purpose the following identity, which is specific to the sinh-Gordon model, is very useful:
$F_{\rm min}(\theta+i\pi)\,F_{\rm min}(\theta)=\dfrac{\sinh\theta}{\sinh\theta+i\sin\pi\gamma}.$ (57)
Using this
$F(\theta+i\pi)=\dfrac{1}{F_{\rm min}(\theta)F_{\rm min}(i\pi)}\,\dfrac{\sinh\theta}{\sinh\theta+i\sin\pi\gamma}.$ (58)
From $S(0)=-1$ one has $\lim_{\theta\to 0}F_{\rm min}(\theta)=0$. In order to obtain a series in powers of $\theta$ we consider
$\partial_{\theta}\log\left(\dfrac{F_{\rm min}(\theta)}{F_{\rm min}(-\theta)}\right)=\partial_{\theta}\log S(\theta).$ (59)
From this one can show
$\lim_{\theta\to 0}F_{\rm min}(\theta)=\dfrac{1}{F_{\rm min}(i\pi)\sin(\pi\gamma)}\left(-i\theta+\dfrac{\theta^{2}}{\sin(\pi\gamma)}+\ldots\right).$ (60)
Thus as expected
$\lim_{\theta\to 0}F(\theta+i\pi)=m^{2}.$ (61)
One can also show for large $|\theta|$:
$\lim_{\theta\to\pm\infty}F_{\rm min}(\theta)\approx 1+\dfrac{i\mathfrak{g}}{4\sinh\theta}.$ (62)
It is also interesting to to study the coupling constant dependence of the basic constant $F_{\rm min}(i\pi)$. From the integral representation (54) one can show
$\log F_{\rm min}(i\pi)=-\gamma+\dfrac{\gamma^{2}\pi^{2}}{8}-\dfrac{\gamma^{3}\pi^{2}}{18}+\dfrac{\gamma^{4}\pi^{4}}{192}+{\cal O}(\gamma^{5}).$ (63)
At self-dual point $b=1$:
$\log F_{\rm min}(i\pi)_{\gamma=1/2}=-\dfrac{2\,{\cal G}}{\pi}+{\rm arccoth}(3),~{}~{}~{}~{}\Longrightarrow~{}~{}F_{\rm min}(i\pi)_{\gamma=1/2}=0.789348.....$ (64)
[…]
Finally let us mention that ${\rho_{\rm vac}}$ can also be determined from a $4$-particle form factor. Define the kernel $G(\theta)$ which appears in the TBA:
$G(\theta)=-i\partial_{\theta}\log S(\theta)=\dfrac{2\cosh\theta\sin\pi\gamma}{\cosh^{2}\theta-\cos^{2}\pi\gamma}.$ (65)
Then one can show LeClairMussardo
$\langle\theta_{2},\theta_{1}|\Theta|\theta_{1},\theta_{2}\rangle=2m^{2}\,G(\theta_{12})\,\cosh(\theta_{12}).$ (66)
Thus one has
$\lim_{\theta_{1}\to\theta_{2}}\,\langle\theta_{2},\theta_{1}|\Theta|\theta_{1},\theta_{2}\rangle=\dfrac{4m^{2}}{\sin(\pi\gamma)}=-32\,{\rho_{\rm vac}}.$ (67)
III.5 Summary
Before turning to higher dimensions D in the next section, it’s useful to summarize the main points of this section. The main new result is the formula (33) which is a prescription for obtaining ${\rho_{\rm vac}}$ from the 2-particle form factor using the bootstrap, and involves a delicate limit where the order of limits matters. Due to the required properties of the crossed form factor, the LHS of this equation is simply $m^{2}$ and this leads to the formula (39). The check of this prescription is that it reproduces the previously known formula from the TBA. Although the formula (39) was derived for diagonal theories, it also applies to non-diagonal scattering, as illustrated for the sine-Gordon model in Mingling . In obtaining (39) it’s important to note that the detailed formula for the 2-particle form factor was not necessary and the final result only depends on S-matrix parameters. We also proposed the principle of particular democracy and showed how it follows from consistency of the S-matrix bootstrap. For the sinh-Gordon model we provided additional analytical details supporting the prescription (33). These results imply that UV divergences in ${\rho_{\rm vac}}$ for free theories can be cured with interactions, where the free field theory limit is $\mathfrak{g}\to 0$.
IV ${\rho_{\rm vac}}$ from the form factor bootstrap in higher dimensions
In this section we will propose a prescription for determining ${\rho_{\rm vac}}$ from the 2-particle form factors of $\Theta$ in D=4 spacetime dimensions. The D=2 result in the last section relied on integrability, which is not available in higher dimensions. In particular for the D=2 theories integrability implies that there is no particle production, the S-matrix factorizes, and this leads to a kind of factorization for $n>2$ particle form factors. Integrability also entails the TBA where our prescription (33) for ${\rho_{\rm vac}}$ can be checked by comparison with the Thermal definition of ${\rho_{\rm vac}}$ proposed in ALCC . None of these integrability features exist for D=4. However bootstrap ideas for massive theories in principle apply to higher dimensions, though in a much more complicated fashion Paulos1 ; Homrich ; Karateev . Moreover, since our prescription is based only on the 2-particle form factor, one does not necessarily need factorization of the S-matrix nor the factorization of the form factor. This opens the possibility that ${\rho_{\rm vac}}$ can be determined from a form factor bootstrap equation based on the 2-particle form factor, and this is the primary focus of this section. In the next subsection we review the basic constraints on the 2-particle form factor for $\Theta$, which only depends on the Mandelstam variable $s$. We then re-write the D=2 result (33) in terms of $s$. The latter result has a natural generalization for ${\rho_{\rm vac}}$ in $D=4$ which we present below.
IV.1 Generalities
For simplicity we assume the theory consists of a single particle of mass $m$ which is its own anti-particle. Throughout this section $F({\bf k}_{1},{\bf k}_{2})$ refers to the 2-particle form factor for the trace of the stress energy tensor $\Theta$ and and $F_{1}$ the 1-particle form factor which is constant:
$F({\bf k}_{1},{\bf k}_{2})\equiv\langle 0|\Theta|{\bf k}_{1},{\bf k}_{2}\rangle,~{}~{}~{}~{}F_{1}=\langle 0|\Theta|{\bf k}\rangle={\rm constant}.$ (68)
Since $\Theta$ is Lorentz invariant $F({\bf k}_{1},{\bf k}_{2})=F(s)$ where $s=(p_{1}+p_{2})^{2}$:
$s=(p_{1}+p_{2})^{2}=2m^{2}+2\left(\omega_{{\bf k}_{1}}\omega_{{\bf k}_{2}}-k_{1}k_{2}\,\cos\phi\right),~{}~{}~{}~{}\widehat{{\bf k}}_{1}\cdot\widehat{{\bf k}}_{2}=\cos\phi.$ (69)
where $\phi$ is the angle between ${\bf k}_{1}$ and ${\bf k}_{2}$. In the center of mass frame, $\cos\phi=-1$ and $|{\bf k}_{1}|=|{\bf k}_{2}|=|{\bf k}|$, thus $s=4m^{2}+2{\bf k}^{2}$.
In any dimension $D$ one still has the Watson equation:
$F(s)=S(s)\,F^{*}(s)+\ldots,$ (70)
where $S(s)$ is the S-matrix for the angular momentum $\ell=0$ partial wave. The additional terms $\ldots$ refer to contributions from thresholds for the production of more than 2 particles, starting at $s>(3m)^{2}$. These are absent for integrable theories in D=2 since there is no particle production, and (70) is equivalent to (26) since $F^{*}(\theta)=F(-\theta)$ and $S^{*}(\theta)S(\theta)=S(-\theta)S(\theta)=1$ due to unitarity. For the remainder of this work we neglect the effect of higher terms in (70). One justification for this is that in $2D$ our prescription for ${\rho_{\rm vac}}$ did not depend on detailed knowledge of the 2-particle form factor. The optical theorem can relate $\Im\,F(s)$ to the 1 particle form factors, and this implies generic poles at $s=m^{2}$. See for instance the arguments in Karateev . However we are interested in bootstrapping $F(s)$ down to the zero-particle form factor $\langle 0|\Theta|0\rangle$, skipping over the intermediate 1-particle form factor, although we will return to it below [^7] The work Karateev focussed mainly on 2 and 1-particle form factors but not the 0-particle form factor considered here..
IV.2 D=2 case in terms of Mandelstam variable $s$
In this sub-section we express our prescription (33) in terms of the Mandelstam variable $s$:
$F(s)\equiv\langle 0|\Theta|{\bf k}_{1},{\bf k}_{2}\rangle,~{}~{}~{}~{}~{}~{}s=(p_{1}+p_{2})^{2}=2m^{2}\left(1+\cosh\theta_{12}\right).$ (71)
Crossing symmetry $p_{1}\to-p_{1}$ implies
$s\to 4m^{2}-s.$ (72)
Thus (39) implies
$\langle{\bf k}|\Theta|{\bf k}\rangle=\lim_{s\to 4m^{2}}F(4m^{2}-s)=m^{2}.$ (73)
The high energy limit of the S-matrix in terms of $s$ is
$\lim_{\theta\to\infty}S(\theta)=1+i\dfrac{\mathfrak{g}}{2\sinh\theta}=1+i\dfrac{\mathfrak{g}\,m^{2}}{\sqrt{s(s-4m^{2})}}.$ (74)
As expected, it has square-root branch cuts originating at $s=0$ and $4m^{2}$. This gives
$\lim_{s\to\infty}(1-S(s))=-i\,\dfrac{\mathfrak{g}\,m^{2}}{s}.$ (75)
Now one has
$\lim_{\theta_{1}\to\infty|\theta_{1}=\theta_{2}}\,s=m^{2}.$ (76)
Thus the prescription (33) reads
$-i\,m^{2}=-i\lim_{s\to\infty|s=4m^{2}}\,F(4m^{2}-s)=\lim_{s\to\infty|s=m^{2}}\left(1-S(s)\right)\,\langle 0|\Theta|0\rangle.$ (77)
This leads to the result in Section III:
$\langle 0|\Theta|0\rangle=\dfrac{m^{2}}{\mathfrak{g}}.$ (78)
IV.3 Prescription for ${\rho_{\rm vac}}$ in D=4 spacetime dimensions
Here one has to work with standard kinematic variables rather than the rapidity, in particular $s$. In the complex $s$-plane, it is known that there are two square-root branch cuts along the real axis, one along the negative axis $s<0$ and another above the 2-particle threshold $s>4m^{2}$. Written in terms of $s$, the equation (77) has a natural generalization to higher dimensions which we propose below. In D=4 dimensions, the states $|{\bf k}\rangle$ have dimension $-1$. Since the operator $\Theta$ has dimension $4$, $\langle 0|\Theta|0\rangle$ has dimension $4$, the two-particle form factor $F(s)$ has dimension $2$ and the 1-particle form factors $F_{1}$ have dimension $3$.
As operators
$S=1+i{\cal T}$ (79)
where ${\cal T}$ contains the interactions. The D=2 result (33) involves $1-S$, thus we consider the diagonal ${\cal T}$ operator matrix element:
${\cal T}(s)\equiv\langle{\bf k}_{1},{\bf k}_{2}|{\cal T}|{\bf k}_{1},{\bf k}_{2}\rangle.$ (80)
Since ${\cal T}$ is dimensionless, ${\cal T}(s)$ has dimension $-4$. In the center of mass frame, the above matrix element is independent of the angle $\phi$ in (69) since the momentum in the $\langle{\rm bra}|$ and $|{\rm ket}\rangle$ are the same. Thus for the matrix element ${\cal T}(s)$ one can consider both the incoming and outgoing particles along the same line in the center of mass frame. The fact that the $\langle{\rm bra}|$ and $|{\rm ket}\rangle$ states in (80) are the same is to be expected: In the Thermal approach, ${\rm Tr}e^{-\beta\,H}$ involves integrating over matrix elements such as in (80). In D=2 dimensions, one-point correlation functions of fields at finite temperature involves such matrix elements LeClairMussardo . These diagonal matrix elements are also central to quantum statistical mechanics formulated completely in terms of the S-matrix Dashen .
[…]
Returning to our problem, based on the correct D=2 result (77) we propose
$\lim_{s\to\infty|s=4m^{2}}\,F(4m^{2}-s)=m^{2}\left(\lim_{s\to\infty|s=m^{2}}{\cal T}(s)\right)\,\langle 0|\Theta|0\rangle.$ (82)
The extra factor of $m^{2}$ on the RHS will serve to define a dimensionless coupling constant $\mathfrak{g}$ below. Now the LHS of the above equation equals $3m^{2}$ by (39). Since ${\cal T}(s)$ has dimension $-4$, let us assume
$\lim_{s\to\infty}{\cal T}(s)=\dfrac{1}{m^{2}}\,\dfrac{\mathfrak{g}}{s}.$ (83)
Then (82) implies
$\langle 0|\Theta|0\rangle=\,\dfrac{3m^{4}}{\mathfrak{g}}~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\rho_{\rm vac}}=\dfrac{3}{4}\,\dfrac{m^{4}}{\mathfrak{g}}.$ (84)
As for the D=2 case, the explicit function $F(s)$ was not needed in order to obtain the above result.
Let us turn now to the 1-point functions, and as before label the particle type as “a”. Then the formula (40), which was based on a cluster decomposition and didn’t rely on integrability, is still expected to hold:
$\lim_{s\to\infty}\,F_{ab}(s)=\dfrac{F_{a}\,F_{b}}{\langle 0|\Theta|0\rangle},~{}~{}~{}~{}~{}F_{a}\equiv\langle 0|\Theta|\theta\rangle_{a}~{}={\rm constant}.$ (85)
The above equation is dimensionally correct, however it can only be valid if $\lim_{s\to\infty}F_{ab}(s)$ is finite. In $2D$ this is ultimately a consequence of (41).
Interestingly, the result (84) is close in spirit to a bound on ${\rho_{\rm vac}}$ which was obtained by very different considerations involving Swampland ideas in connection with electrically charged black holes Montero1 ; Montero2 . There it was proposed that
${\rho_{\rm vac}}<\dfrac{m^{4}}{2e^{2}}$ (86)
where $m$ is the mass of a charged particle, and $\alpha=e^{2}/4\pi$ is the electromagnetic fine structure constant. Whereas the power $m^{4}$ is expected just based on dimensional analysis, the interesting observation is that it is also inversely proportional to an interaction coupling $\mathfrak{g}$. Let us add that for our proposal for ${\rho_{\rm vac}}$ we did not need to assume the particle is electrically charged.
In closing this section, we wish to emphasize that we have not rigorously derived our prescription (82), nor have we have completely justified the high energy limit (83). Rather we based (82) on generalizing the D=2 result (33) in the most natural manner under certain assumptions that we stated. Nevertheless (82) should be viewed as conjectural.
V Potential applications to Beyond the Standard Model physics
Considering the results of the last section, we ask the reader to allow us to take the liberty to speculate on potential implications for physics beyond the Standard Model of particle physics, despite our lack of expertise in its voluminous detailed intricacies. A major open question is the origin of the cosmological constant $\rho_{\Lambda}$, i.e. Dark Energy, which has been measured to be positive and unexplainably small Weinberg ; Martin ; Carroll ; WMAP . There is not yet a consensus on the origins of $\rho_{\Lambda}$ nor its likely resolution. However one prominent idea Weinberg , which is at the origin of the cosmological constant problem, is that it arises from the zero point energy of quantized fields, namely ${\rho_{\rm vac}}$ considered in this paper. [^8] For our perspective on the Cosmological Constant Problem, which is not original, see the short discussion in ALCC and references therein. There actually has not been much serious effort toward a proper calculation of ${\rho_{\rm vac}}$ in the past literature, and one still encounters the incorrect statement that the Standard Model predicts that it is proportional to a cut-off equal to the Planck scale to the $4$-th power. [^9] The latter is the origin of the still often quoted statement that the prediction is off by $120$ orders of magnitude, which is incorrect even in theory. Let us suppose then that ${\rho_{\rm vac}}$ is the only source for $\rho_{\Lambda}$. Then our understanding of ${\rho_{\rm vac}}$ gained in this article has some potential implications for physics beyond the Standard Model.
The Standard Model of particle physics is based on the $SU(3)\otimes SU(2)\otimes U(1)$ Yang-Mills theory, where the $SU(3)$ is QCD. At the electroweak scale, on the order of $M_{\rm Higgs}\approx 125\,{\rm GeV}$, the $SU(2)\otimes U(1)$ is spontaneously broken (SSB) by the Higgs mechanism and this is the standard proposed origin of all particle masses, both for the the quarks and leptons. Simplifying matters, neutrinos only have a left-handed handed helicity, which protects them from obtaining a mass from the Higgs mechanism.
To clarify our reasoning, let us itemize our assumptions, since they certainly require more scrutiny.
Assumption 1.
We assume that the Standard Model is UV complete and in the deep UV it is a CFT. It helps to assume the UV theory is asymptotically free like QCD, which would imply $\lim_{s\to\infty}S(s)=1$, however it’s possible this could be relaxed.
Assumption 2.
Since ${\rho_{\rm vac}}$ is fixed by the deep UV, we assume any Spontaneous Symmetry Breaking (SSB) far below the higher energy scale that determines ${\rho_{\rm vac}}$, such as electroweak symmetry breaking or the de-confinement transition in QCD, does not play a role in determining ${\rho_{\rm vac}}$ since the latter is determined at potentially much higher energies. On the contrary, if such SSB scales were important to ${\rho_{\rm vac}}$, then it is already known that these energy scales are much too high to explain the measured cosmological constant $\rho_{\Lambda}$. To justify this, let us mention that we provided an estimate of ${\rho_{\rm vac}}$ for QCD with 3 massive quarks based on lattice calculations without assuming the de-confinement transition.
Assumption 3.
We assume that in the very deep UV, way above the SSB Higgs scale $M_{\rm Higgs}$, all particles which are thought to obtain their mass from the Higgs mechanism are massless. We mention that our calculation of ${\rho_{\rm vac}}$ for QCD already assumed non-zero quark masses from the Higgs mechanism, and this gave a value of ${\rho_{\rm vac}}$ this is finite and well-defined, but much too high to explain $\rho_{\Lambda}$.
Assumption 4.
In the deep UV, since $\langle 0|\Theta|0\rangle=0$ for a CFT, a non-zero ${\rho_{\rm vac}}$ must arise from at least one non-zero mass particle of mass $m_{\rm z}$ which cannot be explained by the Higgs mechanism. It must then arise from a relevant perturbation of the UV CFT which sets a mass scale. This perturbed theory should not be supersymmetric otherwise ${\rho_{\rm vac}}=0$ by the usual arguments. This does not rule out that the UV CFT is supersymmetric, since the relevant perturbation could break it explicitly. Then the above results of this paper would lead us to propose
${\rho_{\rm vac}}\approx\dfrac{3}{4}\dfrac{m_{\rm z}^{4}}{\mathfrak{g}},$ (87)
where $\mathfrak{g}$ is a dimensionless interaction coupling. For lack of a better name, let us refer to this hypothetical particle as the zero-on, where “zero” refers to the vacuum, or simply the ${\rm zeron}$. This particle is presumed stable, otherwise there would be imaginary parts to poles in $s$ signifying a decay rate which we have not considered.
The measured value of the cosmological constant is  [^10] See for instance WMAP .
$\rho_{\Lambda}\approx 10^{-9}{\rm Joule}/{\rm meter}^{3}\approx(0.003\,{\rm eV})^{4}.$ (88)
Interestingly, there already exists very good candidate for the ${\rm zeron}$, which is the neutrinos. In the Standard Model they only occur with left-handed helicity, thus it is thought that a non-zero neutrino mass cannot arise from the Higgs mechanism. On the other hand, if the neutrino is its own anti-particle, namely a Majorana fermion, since the anti-particle is right-handed, it can pair with the left-handed ones to accommodate a mass term in the action. If this is the case, one expects $\mathfrak{g}$ to be on the order of the fine structure constant $e^{2}/\hbar c\approx 1/137$ for the electro-weak sector. Neglecting overall constants of order $1$, one sees from (88) that if $m_{\rm z}\sim 0.001\,{\rm eV}$ then this could account for the cosmological constant, and $m_{\rm z}$ is close to proposed Majorana neutrino masses Nir .
If the zeron is indeed a massive Majorana fermion, then the formula (84) implies that the Cosmological Constant can be measured from co-linear neutrino-anti-neutrino scattering by measuring $\mathfrak{g}$ based on (83), but only at energies much higher than the electro-weak SSB scale $M_{\rm Higgs}$. The principle of particular democracy implies ${\rho_{\rm vac}}$ can be measured from the scattering of any of such neutrinos. [^11] Incidentally, massive Majorana neutrinos are a leading candidate to explain the matter/anti-matter asymmetry in the current Universe since it can lead to CP violation CPMajorana . Roughly speaking, a massive Dirac fermion can be viewed as two CP conjugate Majorana pairs. It would be quite remarkable if Majorana neutrinos could both explain CP-violation and the Cosmological Constant. Clearly more scrutiny of this idea is worthy of consideration.
VI Conclusions
A main results of this article are the prescription (33) for completing the form factor bootstrap for the trace of the stress-energy tensor, in particular how to obtain the zero particle form factor which determines ${\rho_{\rm vac}}$ from the 2-particle one, and the proposal (82) for its generalization to D=4. This completion requires additional physical input at very high energies which can be determined from the S-matrix. For integrable QFT’s in D=2 spacetime dimensions, our prescription reproduces previously known exact results from the thermodynamic Bethe ansatz. In these even dimensions one finds the simple expression ${\rho_{\rm vac}}\propto m^{\scriptscriptstyle{D}}/\mathfrak{g}$ where $m$ is a physical mass and $\mathfrak{g}$ a dimensionless interaction coup ing which can be inferred from the high energy behavior of the S-matrix. The density ${\rho_{\rm vac}}$ diverges as the interaction $\mathfrak{g}\to 0$ due to known and well-understood UV divergences in free theories. This implies that interactions can potentially cure the known UV divergences in free QFT’s.
With the understanding we obtained from these considerations, we speculated on potential applications to the Standard Model of particle physics. Assuming that the cosmological constant $\rho_{\Lambda}$ comes solely from the vacuum energy density studied in this article, then this would seem to imply the existence of a massive particle that does not obtain its mass from SSB of the electro-weak sector, which we termed the zeron. Massive Majorana neutrinos are a strong candidate for the zeron, since previously proposed neutrino masses approximately have the correct magnitude to account for the astronomically observed $\rho_{\Lambda}$ if $\mathfrak{g}$ is of order 1.
VII Acknowledgements
We wish to thank Peter Lepage, Giuseppe Mussardo and Matthias Neubert for discussions.